OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ \sqrt{9-x ^{2}} }{ x }\]

OpenStudy (anonymous):

\(x=\sin\theta\) I guess.

OpenStudy (anonymous):

asin t correct

OpenStudy (anonymous):

x = 3sint

OpenStudy (anonymous):

Yea. You want it so \(a^2=9\)

OpenStudy (anonymous):

in 3 sin t in for x in the denominator?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yes. for every x in the problem

OpenStudy (anonymous):

\[ dx=3\cos t\;dt \]

OpenStudy (anonymous):

so \[\frac{ \sqrt{9\cos ^{2}\theta} }{ 3\sin \theta }* 3\cos \theta d \theta \]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

then \[\frac{ 9\cos ^{2}\theta }{ 3\sin \theta }d \theta \]

OpenStudy (anonymous):

Nope. it is \[ \frac{ 9|\cos\theta|\cos \theta }{3\sin\theta}d\theta \]

OpenStudy (anonymous):

what are those lines there for?

OpenStudy (anonymous):

Absolute value.

OpenStudy (anonymous):

is it not multiplied?

OpenStudy (anonymous):

Well if \(\cos\theta <0\) then \(|\cos\theta|\cos\theta = -\cos^2\theta\)

OpenStudy (anonymous):

Otherwise it is \(|\cos\theta|\cos\theta = \cos^2\theta\)

OpenStudy (anonymous):

For any definite integral, we would have to split it up for certain intervals of \(\theta\).

OpenStudy (anonymous):

I suppose for simplicity you could just pretend it will always be positive and put a foot note.

OpenStudy (anonymous):

so this is because it is an indefinite integral so cosine could be positive or negative?

OpenStudy (anonymous):

Yeah. It it were a definite integral we would split it up at its roots so we know what intervals it is negative and what intervals it is positive.

OpenStudy (anonymous):

Since it is an indefinite integral, we'd have to make certain assumptions.

OpenStudy (anonymous):

ok that makes sense but now where do we go from here

OpenStudy (anonymous):

Assume it is possible and then integrate.

OpenStudy (anonymous):

positive^

OpenStudy (anonymous):

You basically have \(\tan\theta \cos \theta\)

OpenStudy (anonymous):

You could try parts here I guess

OpenStudy (anonymous):

would that eventually get me to \[3\ln(\frac{ \sqrt{9-x ^{2}}-3 }{ x })+\sqrt{9-x ^{2}}+C\]